OpenStudy (anonymous):

What is the simplified value of the exponential expression 27^1/3 ?

OpenStudy (niyex):

the cube root of 27 \[\sqrt[3]{27}\]

OpenStudy (anonymous):

is that the answer

OpenStudy (g3n3sis08):

\[3\sqrt{3}\]

OpenStudy (g3n3sis08):

The final answer.. if you solve this you get decimals

OpenStudy (chillout):

\[\sqrt[3]{27}=\sqrt[3]{3^3}=3^{3/3}=3\]

OpenStudy (niyex):

genesis08 you are etting it mixed up check out what chillout simplification